Calculate the heat absorbed by 25.0 g of water needed to raise its temperature from 20.0 to 65.0 degrees C. The specific heat of water is 4.18 J/gC
Answer #1

DATE mass of H₂O = 25 gram (m) Specific Heat of water = 4.101/gº T = 2000 ДДТ ЕТ. – Т A . T2 = 65°C AT = 65-20 =45°C We know